OpenStudy (anonymous):

y=-2cos(3x+(pi/2)) Period=2pi/3 Intervals= ? Start and End =? And I must graph it. I am completely Lost Please HELP!

OpenStudy (anonymous):

This is what i did and got no points

OpenStudy (tkhunny):

Can you graph \(y = \cos(x)\)? Start: (0,1) Stop: (2\(\pi\),1) Period: \(2\pi\) Amplitude: 1 Zeros: (\(\pi/2,0\)) and (\(3\pi/2,0\))

OpenStudy (anonymous):

Yes i can

OpenStudy (anonymous):

now will the intervals be [-pi/6,pi/6]

OpenStudy (tkhunny):

How about this one? y=cos(3x)? Start: (0,1) Stop: (2π/3,1) Period: 2π/3 Amplitude: 1 Zeros: (π/6,0) and (π/2,0)